OpenStudy (anonymous):

(Will give medal!) Air is about 78% nitrogen gas (N2) by mass. The molar mass of nitrogen gas is 28.0 g/mol. A 100.0-g sample of air contains how many moles of nitrogen? 2.8 3.6 28 78

OpenStudy (anonymous):

78 % by mass mean. 100 g of air contain 78 g N2. moles of Nitrogen = mass/m.mass=78/28=2.8 moles